A water tank has a hole 3.0 cm^2 in area located 2.0 m below the water surface. How many kilograms of water emerge from the hole per second?



1
Expert's answer
2021-08-10T08:01:31-0400

The mass of water that flows from a hole for time "\Delta t"

"\Delta m=\rho Av\Delta t"

Here

"v=\sqrt{2gH}"

Hence,

"\frac{\Delta m}{\Delta t}=\rho A\sqrt{2gH}\\
=1000*3.0*10^{-4}*\sqrt{2*9.8*2.0}=1.88\:\rm kg/s"